What is the density of CH4 at 27°C and 5 atm pressure
urgent,,please help!!!​

Grade:11

4 Answers

abhay gupta
3 years ago
Explanation:

Here, T=27



C=27+273=300K

P=2.5 atm

The molar mass of(M) CO

2

=44g/mol

From ideal gas equation,

d=

RT

PM

=

0.0821×300

2.5×44

=4.46 g L

−1
abhay gupta
3 years ago
Explanation:

Here, T=27



C=27+273=300K

P=2.5 atm

The molar mass of(M) CO

2

=44g/mol

From ideal gas equation,

d=

RT

PM

=

0.0821×300

2.5×44

=4.46 g L

−1
Sowmiya
2 years ago

PV=nRT

Since V=, where m is the mass of the gas and ρ its  and n=mM, where M is the molar mass of the gas, we can write:

Pmρ=mRTM

We now have to solve for ρ. Dividing both sides by Pm gives:

1ρ=mRTMPm

Which reduces to:

1ρ=RTMP

Reciprocate both sides:

ρ=MPRT.

Here,

M=16.04g/mol

P=1atm

T=273.15K

R=0.0821L atm K−1mol−1.

Inputting:

ρ=16.04⋅10.0821273.15

ρ=0.72

Now we have our number. What about the unit?

We know that density's units are a mass unit divided by a volume unit.

The mass unit here is grams, as the molar mass was expressed in g/mol. The volume unit is liters, as the value of R contained that value. As 1L=1dm3, we can write the density as:

0.72g/dm3

Sowmiya
2 years ago

Here, T=27

 

 

C=27+273=300K

 

P=2.5 atm

 

The molar mass of(M) CO

 

2

 

=44g/mol

 

From ideal gas equation,

 

d=

 

RT

 

PM

 

=

 

0.0821×300

 

2.5×44

 

=4.46 g L

 

−1